Compute the 95 percent interval estimate of ų given ó = 9, n=40 and x= 115

1
Expert's answer
2022-04-14T16:57:49-0400

"95\%CI=(115-1.96\frac{9}{\sqrt40}, 115+1.96\frac{9}{\sqrt40})=(112.21,117.79)."